night-report-2015-03-20

Observer : Heather Campbell
Support team : Southampton (Mat Smith and Régis Cartier)
TAT :Annalisa De Cia and Maria Teresa
Thick clouds in afternoon - hope it clears
Rotator over heated in afternoon, engineers trying to fix. managed to fix it about 22.10

Start of night: Lots of high level thin cloud, better than in the afternoon. wind 2m/s, humidity 60% and rising

End of night: 
Seeing: Measured at start of night, for most of night.

Afternoon calibrations:    EFOSC arcs and specflats and bias

23.02    Sunset

23.30 Spec standard GD71

23.45 PSNJ07361576 Gr11+Gr16

00.20 Change to SOFI
Problem with noise in the SOFI detector, they are looking at it now, we have to wait and see if it can be fixed. Lost ~30mins while being sorted.

01.00  PSNJ07361576 JHK +BG +RG

02.10 Telluric Hip054366 for  PSNJ07361576 

02.20 Telluric Hip029601 for PA15ae

02.30 PS15ae SOFI BG +HK imaging  (seeing around 1.7'' not sure if the imaging is useful)
seeing went up to 2.5'' so stopped, didnt get the J band

04.50 PS15br BG

06.30 Telluric Hip048072  for PS15br

06.40  SNhunt281 BG 2160s +RG 3600s

09.05 Telluric for SNhunt281 Hip060048

09.15  Flux standard: LTT3218 08 41 32.56 -32 56 34.9

09.28    Twilight
09.30 CHANGE TO EFOSC

09.40 SMTJ13455799-3532544  

10.00 Skyflats BR


Comments